Solve using the Factoring method.
8x^2+19x=15

Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(7384)

8x^2+19x=15
8x^2+19x-15 = 0
8x^2+8x-15x-15 = 0
(8x^2+8x)-(15x+15) = 0
8x(x+1)-15(x+1) = 0
(x+1)(8x-15) = 0
x = {-1, 15/8}

Answer by stanbon(75887)

Solve using the Factoring method.
8x^2+19x=15
----
8x^2+19x-15 = 0
(x+3)(8x-5) = 0
---
x = -3 or x = 5/8
====================
Cheers,
Stan H.
==========